If the random variable X takes the values 1, 2, 3 and 4 such that

P(X = 1) = 3P(X = 2) = P(X = 3) = 5P(X = 4), find the probability distribution and

cumulative distribution function of X.



1
Expert's answer
2022-04-08T07:11:01-0400

Let "P(X=1)=p;"

"P(X=2)=\cfrac{p}{3};\\
P(X=3)=p;\\
P(X=4)=\cfrac{p}{5};\\
p+\cfrac{p}{3}+p+\cfrac{p}{5}=1;\\
2p+\cfrac{5p+3p}{30}=1;\\
p=\cfrac{15}{38}."

The probability distribution:

"P(X=1)=\cfrac{15}{38},P(X=2)=\cfrac{5}{38},
\\P(X=3)=\cfrac{15}{38},P(X=4)=\cfrac{3}{38}."



If "x<1, F_X(x)=P(X\le x)=0."

Next, if "1\le x<2, F_X(x)=P(X\le x)=P(X=1)=\cfrac{15}{38}."

Next, if "2\le x<3,"

"F_X(x)=P(X\le x)=P(X=1)+P(X=2)=\cfrac{15}{38}+\cfrac{5}{38}=\cfrac{20}{38}."

Next, if "3\le x<4,"

"F_X(x)=P(X\le x)=\\
=P(X=1)+P(X=2)+P(X=3)=\\
=\cfrac{15}{38}+\cfrac{5}{38}+\cfrac{15}{38}=\cfrac{35}{38}."

Finally, if "x>4,"

"F_X(x)=P(X\le x)=\\
=P(X=1)+P(X=2)+P(X=3)+P(X=4)=\\
=\cfrac{15}{38}+\cfrac{5}{38}+\cfrac{15}{38}+\cfrac{3}{38}=\cfrac{38}{38}=1."


We have the cumulative distribution function of X:

"F_X(x)=\begin{cases}
 0 & & &\text{for } x<1 \\
 \cfrac{15}{38} & & &\text{for } 1\le x<2\\
 \cfrac{20}{38} & & &\text{for } 2\le x<3\\
 \cfrac{35}{38} & & &\text{for } 3\le x<4\\
 1 & & &\text{for } x\ge4.
\end{cases}"